OpenStudy (anonymous):

Write an equation of a line perpendicular to y = x + 4 in slope-intercept form that passes through the point (-2, 6).

OpenStudy (anonymous):

ok so first you need to find the slope of the new line

OpenStudy (anonymous):

wich would be....

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

then you should find the point -2,6 on a graph and count right one unit then down one unit until you hit the y axis, then you have the y intercept

OpenStudy (anonymous):

to find the perpendicular of a line, take the slope's opposite and flip its fraction. The slope here is 1, take the opposite (-1) and flip it (still -1) our new line is y=-x+b plug in the points (-2,6) ---(x,y) in the above line 6= -(-2)+b to get the y intercept y int= 4 final answer is y=-x+4

OpenStudy (anonymous):

ps. the "flip" of the fraction is called its reciprocal

OpenStudy (anonymous):

a. y= 1/4 x - 4 b. y= 1/4x + 6 c. y= -4x - 2 d. y= -4x + 1/4